Hubble Servicing Mission 3A Update - Dec. 22-23 Day 3

Update for Thursday Dec. 23, 05:15 CET

The astronauts will go to sleep around 7:40 CET. The next spacewalk is scheduled for 20:40 CET later this evening. The plan for this second spacewalk on this mission is:

0:00-0:15 Set-up. 0:15-2:05 Replacement of the computer. 2:15-5:15 Replacement of Fine Guidance Sensor 2. 5:15-6:00 Close-out.

Update for Thursday Dec. 23, 04:15 CET

The spacewalk ended officially at 04:10 CET which means that the total duration was 8 hours and 15 minutes. Only one earlier spacewalk, during STS-49, lasted longer - 8 hours 29 minutes.

Update for Thursday Dec. 23, 03:45 CET

The astronauts are now entering the airlock after one of the longest spacewalks in history. The aliveness test for the VIKs turned out positive.

Update for Thursday Dec. 23, 03:30 CET

The crew have now almost finished cleaning the Shuttle bay for tools and equipment. In 20-25 minutes the spacewalk should end.

Update for Thursday Dec. 23, 03:05 CET

The last remaining three VIKs have now been installed.

Update for Thursday Dec. 23, 02:35 CET

Grunsfeld has installed three of the voltage/temperature kits (VIKs), and as a little bonus, Smith has at the same time installed some of the handrailing covers. It has now been decided to install the last three VIKs anyway.

Update for Thursday Dec. 23, 01:55 CET

Smith and Grunsfeld had some problems closing the aft shroud doors. These doors are rather large and a somewhat wobbly. It took a little bit of effort and some time. In fact the team on the First Servicing Mission also had problems closing these doors, but at least this time it went much easier. Due to the time constraints it has been suggested that the team will only change three of the six voltage/temperature kits tonight. The rest will have to be done tomorrow.

Update for Thursday Dec. 23, 01:15 CET

The NICMOS tasks are now completed.

Update for Thursday Dec. 23, 00:38 CET

The crew will now move back to the NICMOS valves and try to force them open.

Update for Thursday Dec. 23, 00:22 CET

It has been confirmed that the third gyro-pair is working. This means that we can say that Hubble has now been repaired! The rest of the mission will mainly be concerned with improvements of Hubble's capabilities.

Update for Thursday Dec. 23, 00:15 CET

The NICMOS task did not succeed, but they have photographed the valves and may return later. Now the team presses on with installation of the temperature/voltage kits which are going to prevent overheating of the batteries onboard. Smith and Grunsfeld have been spacewalking for 4 hours and 20 minutes now.

Update for Thursday Dec. 23, 00:00 CET

The next task is to remove caps on coolant valves on the NICMOS (Near Infrared Camera and Multi-Object Spectrometer) instrument. This is necessary in order to get rid of any left-over nitrogen coolant as a preparation for the installation of a new NICMOS cooler on the next Servicing Mission (SM3B scheduled for 2001). This operation should only take about 15 minutes..

Update for Wednesday Dec. 22, 23:45 CET

The last set of two gyros have now been replaced! The aliveness test is still to be performed.

Update for Wednesday Dec. 22, 23:15 CET

The work is progressing well. STOCC confirmed that the first new set of gyros work. The second set has been installed and are also working perfect.

Update for Wednesday Dec. 22, 22:35 CET

The first two gyros have been replaced. STOCC will now perform an aliveness test to ensure that the two gyros work.

Update for Wednesday Dec. 22, 22:15 CET

The first of the old gyro-sets have been removed and are being stowed.

Update for Wednesday Dec. 22, 21:22 CET

The first set of two gyros has been "turned off" by the crew in the Space Telescope Operations Control Center (STOCC). The new gyro has been unpacked,

Update for Wednesday Dec. 22, 20:55 CET

The set-up is about done now. Next task is the replacement of the gyros.

Update for Wednesday Dec. 22, 20:35 CET

Steve Smith examines some of the handrails on the telescope (which was actually put there for the astronauts to use). He says that the handrails on the Sun side of the telescope have changed colour, but that there are no flaky paint. The paint flakes have been course for concern since small flakes may find their way into sensitive equipment.

Update for Wednesday Dec. 22, 19:54 CET

Smith and Grunsfeld have now switched the power in their spacesuits from "external" (fed by Discovery) to "internal" and are now moving outside. This is the point where the Servicing Mission begins for real.

Update for Wednesday Dec. 22, 19:47 CET

The Shuttle is now changing its position in space to give the astronauts the best protection from the light and radiation from the Sun.

Update for Wednesday Dec. 22, 19:20 CET

The depressurization of the airlock has begun. NASA expresses optimism regarding the possibility to carry out some of the tasks originally scheduled for the fourth spacewalk during the next three spacewalks.

Update for Wednesday Dec. 22, 19:00 CET

Clervoy has just moved the robotic arm in position. It is now ready to support Smith and Grunsfeld in their work. The exchange of the gyros - which should begin in an hours time and last for three hours - is the single most important task on this mission!

Update for Wednesday Dec. 22, 18:50 CET

The schedule for today's spacewalk looks like this:

0:00-1:00 Set-up. 1:00-3:45 All six gyroscopes will be replaced. They are grouped in pairs in the so-called RSUs (Rate Sensor Units). 3:45-4:00 Removal of caps on coolant valves on the NICMOS (Near Infrared Camera and Multi-Object Spectrometer) instrument (preparations for the installation of a new NICMOS cooler on the next Servicing Mission scheduled for 2001). 4:20-5:30 Six voltage/temperature kits (to prevent overheating the batteries) will be installed. 5:30-6:00 Close-out.

We have just received notice from the ground controllers: They have scrutinized the videotapes recorded by the astronauts yesterday, and Hubble does really look like it is in fine shape! Hubble is ready for the first spacewalk.

Update for Wednesday Dec. 22, 18:30 CET

Today's two spacewalkers, Smith and Grunsfeld, have begun breathing pure oxygen before entering the spacesuits which have lower air pressure than the Shuttle. The spacewalk is now scheduled for 19:20 CET, which is more than 1 hour ahead of schedule.

Update for Wednesday Dec. 22, 17:55 CET

The astronauts woke up two hours ago and are now busy preparing for the first spacewalk. The preparations are ahead of schedule and the spacewalk may begin as early as 18:45 CET.